# Dynamical Low-Rank Compression of Neural Networks with Robustness under Adversarial Attacks

### v2

@@ 5.1 Extension to convolutional neural networks
DLRT was extended to convolutional layers in zangrando2023rank by compressing $C$ with a Tucker factorization. Little is gained in compressing the window modes as they are traditionally small. Thus, we only factorize $C$ in the feature modes with output and input feature ranks $r_{O}\ll N_{O}$ and $r_{I}\ll N_{I}$ as
@@
We remark that the extension of Algorithm˜1 to a tensor-valued layer with Tucker factorization only requires to change the truncation step; the SVD is replaced by a truncated Tucker decomposition of $S$. The Tucker bases $U_{i}$ can be augmented in parallel similarly to the matrix case.
Table 2: UCM and Cifar10 benchmark. Clean and adversarial accuracy means and std. devs. of the baseline and regularized low-rank networks for different architectures. We report the low-rank results for unregulated, $\beta=0.0$, and the best performing $\beta$, given in Table˜9. Algorithm˜1 (RobustDLRT)  is able to match or surpass baseline adversarial accuracy values at compression rates of up to $95\%$ in most setups. All runs where RobustDLRT surpasses the uncompressed baseline are highlighted.
@@ 6 Numerical Results
UCM dataset We observe in Table˜2 that Algorithm˜1 can compress the VGG11, VGG16 and ViT-16b networks equally well as the non-regularized low-rank compression and achieves the first goal of high compression values of up to $95\%$ reduction of trainable parameters. Furthermore, the clean accuracy is similar to the non-compressed baseline architecture; thus, we achieve the second goal of (almost) loss-less compression. Noting the adversarial accuracy results under the $\ell^{2}$-FGSM, Jitter, and Mixup attacks with various attack strengths $\epsilon$, we observe that across all tests, the regularized low-rank network of Algorithm˜1 significantly outperforms the non-regularized low-rank network. For the $\ell^{2}$-FGSM attack, our method is able to recover the adversarial accuracy of the baseline network. For Mixup, the regularization almost doubles the baseline accuracy for VGG16. By targeting the condition number of the weights, which gives a bound on the relative growth of the loss w.r.t. the size of the input, we postulate that the large improvement could be attributed to the improved robustness against the scale invariance attack (lin2020nesterov, , Section 3.3) included in Mixup. We refer the reader to Section˜B.1.4 for a precise definition of the Mixup attack featuring scale invariance. However, this hypothesis was not further explored and is delayed to a future work. Finally, we are able to recover half of the lost accuracy in the Jitter attack. Overall, we achieved the third goal of significantly increasing adversarial robustness of the compressed networks. We refer to Table˜9 for the used values of $\beta$ and to Section˜B.3 for extended numerical results for a grid of different $\beta$ values.
@@
Cifar10 dataset We repeat the methodology of the UCM dataset for Cifar10, and observe similar computational results in Table˜2. Furthermore, we compare our method in Table˜4 to several methods of the recent literature, see Section˜3. We compare the adversarial accuracy under the $\ell^{1}$-FGSM attack, see Section˜B.1.2 for details, for consistency with the literature results. We find that our proposed method achieves the highest adversarial validation accuracy for all attack strengths $\epsilon$, even surpassing the baseline adversarial accuracy. Additionally, we find an at least 15% higher compression ratio with RobustDLRT than the second best compression method, CondLR savostianova2023robustlowranktrainingapproximate . A similar experiment for the Projected Gradient Descent (PGD) attack madry2018towards is given in Appendix˜A.
ImageNet1k dataset Finally we repeat the methodology for the ImageNet1k dataset using the ViT-32l vision transformer trained from an ImageNet21k checkpoint, and report the results in Table˜3. The hyperparameters are obtained by an initial sweep and reported in Tables˜8 and 9. RobustDLRT consistently yields higher Top-1/Top-5 accuracy across FGSM and Jitter attacks than DLRT, with especially pronounced gains at larger perturbations (e.g., +9+9 points in Top-1 accuracy under FGSM $\epsilon=0.3$). These trends are consistent with our ViT experiments in Table˜2, demonstrating that adversarial regularization enhances robustness without compromising scalability. We benchmark the training runtime of one ImageNet epoch on an A100 80GB GPU. DLRT requires 26m 07s, while RobustDLRT (with the regularizer) requires 27m 51s, corresponding to an overhead of approximately 3%. This overhead can likely be reduced with further implementation optimizations, indicating that our approach is computationally scalable.
@@
Adversarial Training We evaluate the performance of low-rank training for VGG16 on the UCM dataset using adversarial training. Following goodfellow2014explaining , we use the FGSM attack for different values of $\epsilon$ and train on both 50% clean and attacked images per batch. The results reported in Table˜6 illustrate that RobustDLRT is both compatible with and able to benefit from adversarial training. DLRT without regularization benefits slightly from adversarial training, but exhibits a clear margin to RobustDLRT. RobustDLRT is able to approximately match the non-compressed baseline.
@@ 7 Conclusion
This material is based upon work supported by the Laboratory Directed Research and Development Program of Oak Ridge National Laboratory (ORNL), managed by UT-Battelle, LLC for the U.S. Department of Energy under Contract No. De-AC05-00OR22725.
@@ Appendix A Additional Numerical Results
UCM Dataset
The numerical results for the whitebox FGSM, Jitter, and Mixup adversarial attacks on the VGG16 and VGG11 architectures can be found in Figure˜4, Figure˜5, and Figure˜6. The regularizer confidently increases the adversarial validation accuracy of the networks.
@@
Cifar10 Dataset
We run the same experiment in Table˜4 but with the $\ell^{2}$-PGD attack, which is an iterative version of $\ell^{2}$-FGSM with an random perturbation of the input image as the initial condition. Overall, we see that RobustDLRT is competitive with the other robustness-improving methods when the compression rate is taken into account.
@@ Appendix C Proofs
From (3) and noting $\varsigma_{i}(S^{\top}\!S)=\varsigma_{i}(S)^{2}$, we obtain

### v3

@@ 5.1 Extension to convolutional neural networks
DLRT was extended to convolutional layers in zangrando2023rank by compressing $C$ with a Tucker factorization. Little is gained in compressing the window modes as they are typically small. Thus, we only factorize $C$ in the feature modes with output and input feature ranks $r_{O}\ll N_{O}$ and $r_{I}\ll N_{I}$ as
@@
We remark that the extension of Algorithm˜1 to a tensor-valued layer with Tucker factorization only requires to change the truncation step; the SVD is replaced by a truncated Tucker decomposition of $S$. The Tucker bases $U_{O}$ and $U_{I}$ can be augmented in parallel similarly to the matrix case.
Table 2: UCM and Cifar10 benchmark. Clean and adversarial accuracy means and std. devs. of the baseline and regularized low-rank networks for different architectures. We report the low-rank results for $\beta=0.0$ (DLRT) and the best performing $\beta$ that is given in Table˜9. Algorithm˜1 (RobustDLRT)  is able to match or surpass baseline adversarial accuracy values at compression rates of up to $94\%$ in most setups. All runs where RobustDLRT surpasses the uncompressed baseline are highlighted.
@@ 6 Numerical Results
UCM dataset We observe in Table˜2 that Algorithm˜1 can compress the VGG11, VGG16 and ViT-16b networks equally well as the non-regularized low-rank compression and achieves the first goal of high compression values of up to $94\%$ reduction of trainable parameters. Furthermore, the clean accuracy is similar to the non-compressed baseline architecture; thus, we achieve the second goal of (almost) loss-less compression. Noting the adversarial accuracy results under the $\ell^{2}$-FGSM, Jitter, and Mixup attacks with various attack strengths $\epsilon$, we observe that across all tests, the regularized low-rank network of Algorithm˜1 significantly outperforms the non-regularized low-rank network. For the $\ell^{2}$-FGSM attack, our method is able to recover the adversarial accuracy of the baseline network. For Mixup, the regularization almost doubles the baseline accuracy for VGG16. By targeting the condition number of the weights, which gives a bound on the relative growth of the loss w.r.t. the size of the input, we postulate that the large improvement could be attributed to the improved robustness against the scale invariance attack (lin2020nesterov, , Section 3.3) included in Mixup. We refer the reader to Section˜B.1.4 for a precise definition of the Mixup attack featuring scale invariance. However, this hypothesis was not further explored and is delayed to a future work. Finally, we are able to recover half of the lost accuracy in the Jitter attack. Overall, we achieved the third goal of significantly increasing adversarial robustness of the compressed networks. We refer to Table˜9 for the used values of $\beta$ and Section˜A.1 for extended numerical results.
@@
Cifar10 dataset We repeat the methodology of the UCM dataset for Cifar10, and observe similar computational results in Table˜2. Furthermore, we compare our method in Table˜4 to several methods of the recent literature, see Section˜3. We compare the adversarial accuracy under the $\ell^{1}$-FGSM attack, see Section˜B.1.2 for details, for consistency with the literature results. We find that our proposed method achieves the highest adversarial validation accuracy for all attack strengths $\epsilon$, even surpassing the baseline adversarial accuracy. Additionally, we find an at least 15% higher compression ratio with RobustDLRT than the second best compression method, CondLR savostianova2023robustlowranktrainingapproximate . A similar experiment for the Projected Gradient Descent (PGD) attack madry2018towards is given in Section˜A.2.
ImageNet1k dataset Finally we repeat the methodology for the ImageNet1k dataset, using the ViT-32l vision transformer trained from an ImageNet21k checkpoint, and report the results in Table˜3. The hyperparameters are obtained by an initial sweep and reported in Tables˜8 and 9. RobustDLRT consistently yields higher Top-1/Top-5 accuracy across $\ell^{2}$-FGSM and Jitter attacks than DLRT, with especially pronounced gains at larger perturbations (e.g., +9+9 points in Top-1 accuracy under $\ell^{2}$-FGSM $\epsilon=0.3$). These trends are consistent with our ViT experiments in Table˜2, demonstrating that adversarial regularization enhances robustness without compromising scalability. We benchmark the training runtime of one ImageNet epoch on an A100 80GB GPU. DLRT requires 26m 07s, while RobustDLRT (with the regularizer) requires 27m 51s, corresponding to an overhead of approximately 3%. This overhead can likely be reduced with further implementation optimizations, indicating that our approach is computationally scalable.
@@
Adversarial Training We evaluate the performance of low-rank training for VGG16 on the UCM dataset using adversarial training. Following goodfellow2014explaining , we use the $\ell^{2}$-FGSM attack for different values of $\epsilon$ and train on both 50% clean and attacked images per batch. The results reported in Table˜6 illustrate that RobustDLRT is both compatible with and able to benefit from adversarial training. DLRT without regularization benefits from adversarial training, but exhibits a clear margin to RobustDLRT. Additionally, RobustDLRT is able to approximately match the non-compressed baseline.
@@ 7 Conclusion
This material is based upon work supported by the Artificial Intelligence Initiative of the Laboratory Directed Research and Development Program of Oak Ridge National Laboratory (ORNL), managed by UT-Battelle, LLC for the U.S. Department of Energy under Contract No. De-AC05-00OR22725.
@@ Appendix A Additional Numerical Results
A.1 UCM Dataset
The numerical results for the whitebox $\ell^{2}$-FGSM, Jitter, and Mixup adversarial attacks on the VGG16 and VGG11 architectures can be found in Figure˜4, Figure˜5, and Figure˜6. The regularizer confidently increases the adversarial validation accuracy of the networks.
@@ A.1 UCM Dataset
A.2 Cifar10 Dataset
We run the same experiment in Table˜4 but with the $\ell^{2}$-PGD attack, which is an iterative version of $\ell^{2}$-FGSM with an random perturbation of the input image as the initial condition [30]. Overall, we see that RobustDLRT is competitive with the other robustness-improving methods when the compression rate is taken into account.
@@ Appendix C Proofs
From (26) there holds